OpenStudy (anonymous):

The electrical resistance R of a wire varies directly as the length l of the wire and inversely as the square of its diameter d. If 20 meters of wire of diameter 1.5 mm has a resistance of 12 ohms, what is the resistance of 20 meters of the same wire if the diameter is increased to 5 mm?

OpenStudy (ybarrap):

First solve for k, then find R with the new diameter: $$ R=\cfrac{kl}{d^2}\\ 12=\cfrac{k\times20}{(.0015)^2}\\ \implies\\ k=\cfrac{12\times(.0015)^2}{20}\\ R=\cfrac{20k}{(.005)^2} $$

OpenStudy (anonymous):

Im really confused. Here on the answers they give me to choose from: 1.08 5 42 85.3

OpenStudy (ybarrap):

Do you understand the 1st equation why we have l and k at the top and d at the bottom? k is just a constant of proportionality when given that something is proportional to another it only varies by a factor. I'm calling that factor k. Also, I've converted from mm to meters. 1.5 mm = .0015 m and 5 mm = .005 m

OpenStudy (ybarrap):

Were you able to calculate the factor k?

OpenStudy (ybarrap):

If you run down the calculations I've given above, one of your options appears quite easily. Good luck.

OpenStudy (anonymous):

I got 1.35E-6

OpenStudy (anonymous):

OMG I got 1.08 which is an answer to choose from. Thanks!!! (is this correct?)

OpenStudy (ybarrap):

yw